OpenStudy (anonymous):

A rectangular tank has a base that is 16 inches by 5 inches and a height of 18 inches. If the tank is half full of water, by how many inches will the water level rise if 260 cubic inches of water are poured into it?

OpenStudy (anonymous):

tips?

OpenStudy (anonymous):

first calculate the volume=16*5*18...then calculate the half volume...and 260 into that and substarct from the total volume

OpenStudy (anonymous):

k

OpenStudy (anonymous):

the total volume is 1,440

OpenStudy (anonymous):

half of the volume is 720

OpenStudy (anonymous):

so

OpenStudy (anonymous):

wait...

OpenStudy (anonymous):

so

OpenStudy (anonymous):

here's the deal

OpenStudy (anonymous):

the whole volume is 1,440

OpenStudy (anonymous):

and half the volume is 720

OpenStudy (anonymous):

now, if I add 260 cubic inches of water...

OpenStudy (anonymous):

how high it will rise in inches?

OpenStudy (anonymous):

tips?

OpenStudy (anonymous):

what i said add halkf volume to 260 what it becomes

OpenStudy (anonymous):

well..

OpenStudy (anonymous):

here's the thing

OpenStudy (anonymous):

the volume only rises 13/4 inches

OpenStudy (anonymous):

but why?

OpenStudy (phi):

the volume of 260 cc is base * height base is 5*16 height is to be found 80*h= 260 solve for h

OpenStudy (anonymous):

260 cubic inches were added to a body of 720 cubic inches, right?

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

so

OpenStudy (anonymous):

3.25

OpenStudy (anonymous):

ahh

OpenStudy (anonymous):

rise is the key word

OpenStudy (phi):

3.25 is also 13/4 inches

OpenStudy (anonymous):

rise means in height

OpenStudy (anonymous):

ahh...

OpenStudy (phi):

the half full part does not affect anything... unless the 260 cc caused the tank to overflow. Not a problem here.

OpenStudy (aripotta):

 12.25 - 9 = 3.25. the water raised 3.25 inches.